OpenStudy (anonymous):

need help with x/2-x/8=3/4

hero (hero):

Factor out the X

OpenStudy (anonymous):

ok will it be 2-8=3/4

hero (hero):

That way you'll have just x(1/2-1/8) = 3/4

OpenStudy (anonymous):

ok can i multiply tp get 1/16=3/4

hero (hero):

The X doesn't disappear. And also you have to properly subtract 1/2-1/8

OpenStudy (anonymous):

I dont understand.